Use Lagrange multipliers to find the point (a,b) on the graph of y=e^8x, where the value ab is as small as possible.

P= ?
1
Expert's answer
2020-04-07T13:21:53-0400

Consider the objective function "f(x,y)=xy"

Let the constraint function be "g(x,y)=y-e^{8x}"

Apply Lagrange Multiplier as follows:


"\nabla f(x,y)=\lambda \nabla g(x,y)"

So, find the gradients both sides and equate as,


"<y,x>=\lambda <-8e^{8x},1>"

Equate both sides to obtain,


"y=-\lambda8e^{8x}""x=\lambda"

Substitute "\lambda=x" into equation "y=-\lambda8e^{8x}" to obtain "y=-8xe^{8x}"


Now, substitute "-8xe^{8x}" for "y" into constraint equation "y-e^{8x}=0" and solve for "x" as,



"-8xe^{8x}-e^{8x}=0"

Solve the equation as follows:


"e^{8x}(-8x-1)=0"


"-8x-1=0,e^{8x}\ne0"


"x=-\frac{1}{8}"

Next, plug "x=-\frac{1}{8}" into relation "y=-8xe^{8x}" to find the value of "y" as,


"y=-8(-\frac{1}{8})e^{8(-\frac{1}{8})}=e^{-1}"



Therefore, the point "(a,b)" on the graph of "y=e^{8x}" is "(-\frac{1}{8},\frac{1}{e})"